OpenStudy (anonymous):

Please Help.... How many atoms are in 1 g C? Answer in units of atoms

OpenStudy (chmvijay):

can you tell me atomic weight of Carbon

OpenStudy (anonymous):

12 @chmvijay

OpenStudy (chmvijay):

1 mole of Carbon mean how much gram of carbon

OpenStudy (chmvijay):

1Mole is nothing but is equal to atomic weight of element or molecular weight of molecule can you now answer the above

OpenStudy (anonymous):

so its 12 atoms? @chmvijay

OpenStudy (chmvijay):

no no its 12 gram

OpenStudy (chmvijay):

1 mole of C = 12 g of C = 6.023*10^23 atoms of C

OpenStudy (chmvijay):

now 12 gram of carbon contains 6.023*10^23 atoms of carbon then 1 gram of carbon contains= 6.023*10^23 *1 /12 = can u solve it

OpenStudy (anonymous):

No=( I dont have a scientific calculator @chmvijay

OpenStudy (chmvijay):

you just divide 6.023/12

mathslover (mathslover):

Approx. it @JazminZamarripa . Try to do calculation on your own.

OpenStudy (chmvijay):

you just divide 6.023/12 and keep the 10^23 like that only

OpenStudy (anonymous):

its 0.50 @chmvijay

mathslover (mathslover):

Yeah!

OpenStudy (chmvijay):

yaaa fine 0.5*10^23 this is the answer or 5*10^22 also u can write OK

OpenStudy (anonymous):

ok thank you @chmvijay

mathslover (mathslover):

Don't forget to medal others @jazminzamarripa .

mathslover (mathslover):

Award @chmvijay by giving a medal!

mathslover (mathslover):

Good!